### v2

# DeepEN: A Deep Reinforcement Learning Framework for Personalized Enteral Nutrition in Critical Care

Abstract
Objective: Current ICU enteral feeding remains sub-optimal due to limited personalization and ongoing uncertainty about appropriate calorie, protein, and fluid targets—particularly in the context of rapidly changing metabolic demands and heterogeneous responses to therapeutic interventions. This study introduces DeepEN, a novel reinforcement learning (RL)–based framework designed to dynamically personalize enteral nutrition (EN) dosing for critically ill patients using electronic health record data.
Methods: DeepEN was trained on data from over 11,000 ICU patients in the MIMIC-IV database to generate 4-hourly, patient-specific targets for caloric, protein, and fluid intake. The model’s state space integrates demographics, comorbidities, vital signs, laboratory measurements, and recent interventions considered relevant to nutritional management. The reward function was designed with domain expertise to balance short-term physiological and nutrition-related goals with long-term survival outcomes, reflecting real-world clinical priorities. The framework employs a dueling double deep Q-network with Conservative Q-Learning regularization to ensure safe and reliable policy learning from retrospective data. Model performance was benchmarked against both clinician-derived and guideline-based policies.
Results: DeepEN outperformed both clinician and guideline-based policies, achieving a 3.7 ± 0.17 percentage-point absolute reduction in estimated mortality compared with the clinician policy (18.8% vs 22.5%) and higher expected returns relative to the gold-standard guideline policy (11.89 vs 8.11). Control of key nutritional biomarkers was also improved under the learned policy. U-shaped associations were observed between deviations from clinician dosing and mortality for caloric, protein, and fluid recommendations, suggesting that the learned policy aligns with clinician actions associated with higher survival while diverging from suboptimal behaviors.
Conclusion: DeepEN demonstrates the feasibility and potential of conservative offline RL for safe, individualized enteral nutrition therapy in critical care. These findings suggest that data-driven personalization may enhance patient outcomes beyond traditional guideline- or heuristic-based approaches.
keywords:
Reinforcement learning , critical care , enteral nutrition , personalized medicine
Jointly supervised this work.
*Corresponding author: ephfm@nus.edu.sg
1 Introduction
Enteral nutrition (EN), also known as tube feeding or enteral feeding, refers to the supply of nutrients such as carbohydrates, proteins, and fluids directly through the gastrointestinal tract. EN is typically initiated within 48 hours of Intensive Care Unit (ICU) admission for patients who are unable to meet their nutritional requirements through oral intake [adeyinka2018enteric]. This can be due to increased metabolic demand or impairments in swallowing, nutrient absorption, or appetite regulation. Such conditions include mechanical ventilation, severe head injury, altered mental status, and gastrointestinal dysfunction [preiser2021guide]. The provision of adequate calories and protein is essential for such patients, but it remains clinically challenging.
Evidence based guidelines such as from the American Society for Parenteral and Enteral Nutrition (ASPEN) provide weight and condition specific daily targets for calories and protein intake, considering factors such as obesity, renal replacement therapy, and hypermetabolic states [compher2022guidelines]. However, these guidelines would continue to evolve as new clinical evidence emerge. Recent randomized controlled trials comparing aggressive verses conservative nutritional strategies in critically ill patients have reported conflicting results, ranging from improved outcomes to even potential harm [compher2022guidelines, bels2024effect, heyland2023effect]. These mixed findings underscore persistent uncertainties around the optimal dose and timing of enteral nutrition in the ICU setting.
Moreover, there is a well-documented gap between guidelines and clinical practice. Surveys of ICU physicians and nurses reveal that lack of familiarity with both international and local nutrition guidelines, coupled with perceived gaps in the supporting evidence base, significantly hinders the consistent implementation of recommended enteral nutrition practices [mirhosiny2021physicians, friesecke2014improvement, jarden2015practice, ellis2015improving] . Furthermore, nutritional decisions in the ICU are typically guided by heuristic-based practices or one-size-fits-all protocols, which fail to account for the dynamic and patient-specific trajectories of critically ill patients. Moreover factors such as evolving organ function, variable tolerance to feeding, competing interventions (e.g., vasopressors, sedation, dialysis) and rapidly varying metabolic demands make nutritional needs highly individualized and temporally variable. As a result, intervention inconsistencies such as underfeeding, overfeeding and delayed initiation remain common, contributing to increased infection risk, prolonged mechanical ventilation, muscle wasting, and mortality [ramaswamy2024nine].
Consequently, there is a pressing need address these limitations using adaptive, data-driven decision support systems that can effectively recommend personalized nutrition therapy in real time. Reinforcement learning (RL), which is a class of machine learning algorithms that learn optimal decision-making policies through interaction with complex environments, offers a promising framework for this task. RL models can learn from retrospective clinical data to recommend patient-specific, time-dependent feeding strategies that optimize long-term outcomes such as survival and recovery. Unlike static and population-wide guidelines, RL approaches can continuously adapt to a patient’s evolving physiological state, balance competing clinical goals, and offer individualized recommendation at each time point. While RL has shown promise in healthcare settings such as in sepsis [raghu2017deep, tan2024advancing] and mechanical ventilation management [peine2021development], to our knowledge, it has yet to be utilized in the domain of enteral nutrition in critical care.
Our main contributions are as follows:
  • 1.
    We introduce DeepEN, the first RL-based framework for EN management in critical care. DeepEN generates dynamic, patient-specific recommendations for caloric, protein, and fluid targets, optimizing long term clinical outcomes. DeepEN conditions its decisions on a comprehensive list of patient features including diagnoses, labs, vitals, and past treatments to make personalized recommendations. The solution uses conservative RL techniques to ensure clinically safe and plausible recommendations when learning from retrospective data.
  • 2.
    We conduct a comprehensive evaluation of DeepEN’s performance against current clinical practice and established clinical guidelines. We evaluate the learned policies using well-established qualitative and quantitative off-policy evaluation methods, and showcase DeepEN’s ability to support more consistent, individualized, and outcome-driven nutritional care in critical care.
  • 3.
    We introduce a domain-specific reward function tailored to the context of critical care nutrition. This reward function balances both long- and short- term clinical goals by integrating intermediate signals with terminal outcomes, allowing more clinically aligned policy optimization in the complex, high-stakes setting of the ICU.
2 Background and Related Work
2.1 Reinforcement Learning
Reinforcement Learning (RL) is a framework used to optimize sequential decision-making processes. It is typically modeled as a Markov Decision Process (MDP), which is defined by the tuple $\left(\mathcal{S},\mathcal{A},r,P,\gamma\right)$, where $\mathcal{S}$ is the set of states, $\mathcal{A}$ is the set of actions, $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ is the reward function, $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$ is the transition probability function, and $\gamma\in(0,1)$ is the discount factor.
At each discrete time step $t$, an agent observes its current state $s_{t}\in\mathcal{S}$, selects an action $a_{t}\in\mathcal{A}$, and transitions to a new state $s_{t+1}$ according to the transition dynamics $P(s_{t+1}\mid s_{t},a_{t})$, while receiving a reward $r_{t}=r(s_{t},a_{t})$.
The overarching goal in reinforcement learning is to discover a policy $\pi:\mathcal{S}\to\mathcal{A}$ that maximizes the agent’s expected cumulative discounted return, often defined as the sum of future rewards: $\sum_{t=0}^{T}\gamma^{t}r_{t},$ where $T$ is the time horizon.
2.2 Q-learning and Offline RL
Q-learning [watkins1992q] is one of the foundational algorithms in reinforcement learning (RL) and is among the most commonly used approaches in healthcare RL. [yu2021reinforcement] The method aims to learn the value of taking an action $a$ in a given state $s$, known as the Q-value $Q(s,a)$. At each time step $t$, the agent updates this estimate based on the reward $r_{t}$ received and the expected future return, using the Bellman update rule:
$Q(s_{t},a_{t})\leftarrow Q(s_{t},a_{t})+\eta\left(r_{t}+\gamma\max_{a^{\prime}}Q(s_{t+1},a^{\prime})-Q(s_{t},a_{t})\right),$
where $\eta$ is the learning rate and $\gamma\in(0,1)$ is the discount factor applied to future rewards.
To handle large or continuous state spaces, Deep Q-Networks (DQN) [mnih2015human] approximate the Q-function using a neural network. While this leverages the expressive power of neural networks, it also introduces sensitivity to the distribution of training data. In particular, the presence of the maximization operator $\max_{a^{\prime}}Q(s_{t+1},a^{\prime})$ in the update rule can cause upward bias in the target, especially when Q-values are based on noisy or overfitted estimates, leading to overestimation errors.
This issue becomes more profound in offline (or batch) reinforcement learning, where the agent is limited to learning from a fixed dataset $\mathcal{D}$ collected by a behavior policy [levine2020offline]. In healthcare RL, offline RL is the standard approach since direct interaction with the environment (i.e., directly acting on real patients) is unsafe and ethically impermissible. Without access to environmental interaction (as in online RL), the algorithm must generalize from the static dataset. Consequently, state-action pairs that are rare or missing in $\mathcal{D}$ may result in out-of-distribution (OOD) estimates, exacerbating overestimation error. In high-stakes domains like healthcare, such inaccurate Q-values can lead to unsafe or clinically inappropriate treatment recommendations if not properly mitigated.
2.3 Addressing overestimation bias with D3QN and CQL in offline RL
To mitigate the overestimation bias inherent in standard DQN, the Dueling Double DQN (D3QN) [van2016deep, wang2016dueling, hessel2018rainbow] algorithm was proposed as an enhancement. D3QN incorporates two key enhancements. First, it employs the Double Q-learning approach to decouple action selection and evaluation. The action is selected using the online network as $a^{\prime}=\arg\max_{a}Q(s_{t+1},a;\theta)$, and evaluates using a separate target network as $Q(s_{t+1},a^{\prime};\theta^{-})$. Second, D3QN incorporates a dueling network architecture that separately estimates the state-value function and the advantage function. This structure allows the agent to more effectively learn which states are valuable independently of the chosen actions, improving learning efficiency and stability. While these augmentations reduce the overestimation bias caused by the maximization operator over noisy Q-estimates, D3QN still does not fully resolve challenges in offline reinforcement learning, where distributional shift and insufficient coverage of state-action pairs remain significant limitations.
To further combat overestimation in offline reinforcement learning, Conservative Q-Learning (CQL) [kumar2020conservative] was proposed. CQL addresses the risks posed by out-of-distribution (OOD) state-action pairs by explicitly penalizing overestimated Q-values for unseen or rarely seen actions. Specifically, CQL augments the Q-learning error with the term $\mathbb{E}_{s_{t}\sim\mathcal{D},a_{t}\sim\mathcal{A}}[Q(s_{t},a_{t})]$, which penalizes large Q-values for all possible actions at a given state, including those not sampled in the dataset. This discourages the agent from overvaluing out-of-distribution actions. In addition, CQL also includes the negative expectation $\mathbb{E}_{(s_{t},a_{t})\sim\mathcal{D}}[Q(s_{t},a_{t})],$ which promotes higher Q-values for the actions that are actually observed in the dataset. Together, these terms ensure that the learned Q-function remains pessimistic about unsupported actions while maintaining high values for well-supported actions, implicitly steering the learned policy towards in-distribution (’clinician-like’) behavior. We combine both techniques of D3QN and CQL to avoid unsafe policy recommendations driven by overestimated Q-values in underrepresented areas of the state-action space.
2.4 RL for Enteral Nutrition Management
In the context of clinical decision support, RL has been widely explored for applications such as fluid and vasopressor dosing in sepsis patients, [raghu2017deep, tan2024advancing], mechanical ventilation management, [peine2021development], and sedation titration [yu2021reinforcement]. However, to the best of our knowledge, RL has not yet been applied to the management of enteral nutrition (EN) in the ICU. Existing AI efforts in this domain have largely focused on predictive tools for malnutrition screening [besculides2023implementing, raphaeli2023using] or detecting intolerance to EN [hu2022development, chen2023development], rather than personalized nutritional support. Developing an RL-based nutrition management system presents both novel opportunities and unique challenges. Unlike pharmacologic or mechanical interventions (e.g., vasopressors or ventilator settings) that typically produce immediate and measurable physiological responses, nutritional interventions yield more gradual and diffuse effects, and often confounded by concurrent therapies and patient heterogeneity. This complicates the construction of a reward function that meaningfully encodes clinical outcomes such as mortality, recovery, or nutritional adequacy.
Furthermore, as EN remains a largely unexplored use case for RL, the design of the state representation and reward function requires careful consideration. These components must be constructed by leveraging both clinical expertise and empirical analysis to ensure the agent can learn relevant dynamics and make safe, effective recommendations. Our work represents a first step toward leveraging offline RL for safe and effective personalization of nutritional support in critical care.
3 Methodology
This section details our methodology, including RL problem formulation, algorithm design, cohort construction and experiment setup. Figure 1 illustrates an overview of the solution pipeline.
Figure 1: Overview of the DeepEN solution pipeline.
3.1 RL Problem Definition
3.1.1 States
Our state space (S) comprises 102 variables consisting 63 base variables outlined in table 1, and an additional 39 ‘rate-of-change’ variables.1 The rate-of-change variables are the average rate of change calculated across the previous three time-steps for 39 chosen base variables. The chosen variables have dynamics that reflect clinically important trends. These were included to capture additional contextual information on the patient’s health trajectory and disease progression.
Table 1: State Space Variables
Category Variables
Demographics (4) Age, Gender, Weight*, ICU readmission
Diagnoses and Comorbidities (6) Burns, CKD, Diabetes, Sepsis, Trauma, Elixhauser score
Vitals (13)* HR, SBP, MBP, DBP, Resp. rate, Temperature, PaCO2, PaO2, PF ratio, SpO2, SOFA, GCS, Shock index
Labs (24)* Albumin, pH, Calcium, Glucose, Hemoglobin, Magnesium, WBC, Creatinine, Bicarbonate, Sodium, Lactate, Chloride, Platelets, Potassium, PTT, PT, AST, ALT, BUN, INR, Ionised calcium, Total bilirubin, Base excess, Phosphate
Feeding Related (5) Previous calories, Previous protein, Previous water, Cumulative calories, Cumulative protein
Treatments and Interventions (8) Mechanical ventilation, FiO2, CRRT, IV fluids, Vasopressor dose, Propofol dose, Insulin dose, 24-h cumulative insulin dose
Others (3) Urine output (4-hourly)*, Total output, Time since EN initiation
* Variables whose average rate of change was also included.
Abbreviations: CKD = chronic kidney disease; DBP = diastolic blood pressure; GCS = Glasgow Coma Scale; HR = heart rate; ICU = intensive care unit; MBP = mean blood pressure; PTT = partial thromboplastin time; Resp. = respiratory rate; SBP = systolic blood pressure; WBC = white blood cell.
3.1.2 Actions
The action space is defined using three main components of enteral nutrition: calories, protein and water. For each patient and time window, we consider the weight-adjusted quantities of these component, each discretized into four levels based on empirical quantiles. In line with clinical guidelines, we define total energy and fluid intake as the sum of EN and non-EN sources (e.g., IV fluids and propofol). Although non-EN sources are included for completeness, EN make up the vast majority of these nutritional components.
The action space A is the Cartesian product of the set of these three components, resulting in a theoretical total of 64 ($4\times 4\times 4$) possible action combinations. However, due to inherent correlation among calories, protein and water administration, only 51 of the combinations are observed in the dataset. Hence these combinations are considered as the functional action space. Note that the quantities of all three components are strictly positive, as we focus only on periods where enteral nutrition was actively administered. Further details on the action composition and discretization process is detailed in the Appendix (Section 2: Action Information).
3.1.3 Rewards
We define a composite reward function that integrates both intermediate physiological indicators and long-term patient outcomes. The formulation is inspired by well-established reward structures used in healthcare RL research [raghu2017deep], and tailored to the clinical context of enteral nutrition.
Let $R_{t}\in\mathbb{R}$ denote the scalar reward assigned at time step $t$. For each patient trajectory, rewards are defined as:
$R_{t}=\begin{cases}R_{\text{term}}(m),&\text{if }d_{t}=1\\ R_{\text{im}}(t),&\text{if }d_{t}=0\end{cases}$
where $d_{t}\in\{0,1\}$ is an indicator of whether the current state is terminal, $m\in\{0,1\}$ denotes ICU mortality (1 for death, 0 for survival), $s_{t}$ and $s_{t+1}$ are the current and next state vectors, $R_{\text{term}}(m)$ is the terminal reward assigned based on mortality outcome, and $R_{\text{im}}(t)$ are the intermediate reward components based on physiological and biomarker changes, respectively.
Terminal Reward ($R_{\text{term}}(m)$):
The terminal reward is a scalar assigned at the final time step and is defined as:
$R_{\text{term}}(m)=\begin{cases}+r_{T},&\text{if }m=0\\ -r_{T},&\text{if }m=1\end{cases}$
Consistent with established reward formulations in prior work in RL-based dynamic treatment regimes [raghu2017deep, tan2024advancing], we set $r_{T}=15$, which preserves the dominance of the terminal outcome while remaining compatible with the scale of intermediate shaping signals.
Intermediate Rewards ($R_{\text{im}}(m)$):
To mitigate the challenge of sparse rewards and enable efficient credit assignment, we introduce intermediate rewards at each non-terminal timestep. This design facilitates learning by providing frequent and interpretable signals about the clinical trajectory as follows:
$R_{\text{im}}(t)=R_{\text{phys}}(s_{t},s_{t+1})+R_{\text{bio}}(s_{t},s_{t+1})$
By shaping rewards using short-term physiological responses and penalizing biomarker deviations, the agent is encouraged to learn policies consistent with safe, clinically appropriate trajectories, even when terminal outcomes are not immediately observable.
Intermediate Physiological Reward ($R_{\text{phys})}$: The physiological component helps guide clinical improvement in organ function and perfusion. Based on prior work in critical care RL, it is designed to encourage immediate improvement in SOFA score and blood lactate as follows:
$R_{\mathrm{phys}}={}$ $-c_{0}\mathbf{1}\{\,\Delta\mathrm{SOFA}=0\ \land\ \mathrm{SOFA}_{t}>0\,\}$ (stagnation) (1)
$-c_{1}(\mathrm{SOFA}_{t+1}-\mathrm{SOFA}_{t})$ (SOFA change)
$-c_{2}\,\tanh(\mathrm{Lactate}_{t+1}-\mathrm{Lactate}_{t})$ (lactate penalty)
where $c_{0}=0.025$, $c_{1}=0.125$, and $c_{2}=2.0$ are fixed scalar weights chosen based on prior validation. The formulation rewards reductions in severity of organ dysfunction and lactate accumulation while penalizing stagnation or worsening trends.
Intermediate Biomarker Deviation Reward ($R_{\text{bio}}$): We further incorporate two biomarkers — blood glucose and serum phosphate — due to their nutritional relevance and responsiveness to carbohydrate and protein intake [gosmanov2013management, uribarri2007phosphorus, kalantar2010understanding, ikizler2013prevention]. These biomarkers serve as strong signals indicative of feeding tolerance and general safety-related clinical outcomes [mehanna2008refeeding, nguyen2007relationship]. Interventions such as insulin, which can directly alter serum glucose levels, were included in the state space to account for their confounding effects on these biomarkers. The composite biomarker reward is defined as:
$R_{\text{bio}}=\lambda_{g}\,R_{g}+\lambda_{p}\,R_{p}$
where $R_{g}$ and $R_{p}$ are individual reward terms for glucose and phosphate, respectively. Both weighting coefficients were set to $\lambda_{g}=\lambda_{p}=1$ in the final policy to balance the biomarkers and constrain the reward’s magnitude. This was done to reduce the risk of short-term biomarker fluctuations overshadowing critical long-term outcomes while maintaining fidelity to domain priorities.
Each component reward is calculated using a smooth sigmoid shaping function and a bonus term for improvement:
$R_{x}=f_{x}(x_{t+1})+\varepsilon\,\bigl(\delta_{x}^{\text{pre}}-\delta_{x}^{\text{post}}\bigr)_{+},$
where $(\cdot)_{+}$ denotes the positive part operator, ensuring the bonus is non-negative. The term $\varepsilon$ (set to 0.2) weights the bonus for decreasing deviation, and
$\delta_{x}=\frac{\max\!\bigl(0,\,x-x_{\max}\bigr)+\max\!\bigl(0,\,x_{\min}-x\bigr)}{x_{\max}-x_{\min}}$
is the normalized deviation from the clinical target range.
The shaping function $f_{x}$ is implemented as a difference of two logistic curves centred at the lower and upper bounds of the target range:
$f_{x}(x)=\frac{2}{1+\exp\!\bigl(-(x-x_{\min})\bigr)}-\frac{2}{1+\exp\!\bigl(-(x-x_{\max})\bigr)}$
This function yields a smooth plateau when $x$ lies inside $x_{\min},x_{\max}]$ and decays gradually outside the interval, encouraging maintenance of the biomarker within the desired range rather than penalising small fluctuations.
For glucose, the target range is $[140,180]$ mg/dL [american202416], and for phosphate, the target range is $[2.5,4.5]$ mg/dL [geerse2010treatment]. This formulation rewards both the absolute proximity of each biomarker to its target and any improvement over time.
3.2 Experimental Setup
3.2.1 Data Extraction and Preprocessing
We extracted a cohort of enteral feeding patients from the MIMIC-IV database [johnson2023mimic] which contains data for over 60,000 ICU patients from the Beth Israel Deaconess Medical Center (Boston, MA, USA) from between 2008 to 2019. Inclusion criteria contained following conditions: 1) patient is >>18 years of age, 2) has at least 12 hours of enteral feeding data, and 3) has at least one recorded weight measurement. In total, data for 11,378 patients were included. For each patient, we collect data on demographics, vital signs, lab-values, fluids, specific diagnoses, enteral feeding data, and other relevant variables detailed in section 3.2. We only include data from periods where EN is administered, and did not consider periods where EN is absent or interrupted, nor did we consider data from parenteral feeding. In line with clinical evidence, we use enteral feeding data only from the first 10 days post-ICU admission to focus on the acute phase of critical illness (PCI) [iwashyna2015towards]. Data for calories, protein, and water intake were extracted from MIMIC-IV’s inputevents table. We aggregated the patient trajectories into four-hourly windows using mean or sum as appropriate, and fill any missing values using linear interpolation. More details on the patient cohort can be found in the Appendix (see Section 1: Cohort Information).
3.2.2 Baselines
To benchmark the performance of our reinforcement learning (RL) policy, we compare it against four distinct baselines: a random dosage policy, a clinician policy, a behavior cloning (BC) policy, and a expert guideline-based policy.
  1. 1.
    Random Dosage Policy: The random policy selects actions uniformly at random from the discrete action space. Formally, it is defined as $\pi(a)=\frac{1}{M}$, where $M$ is the total number of possible actions. This baseline serves as a critical sanity check, establishing a worst-case scenario or lower performance bound. It allows us to assess whether the learned policy outperforms naive, uninformative behavior.
  2. 2.
    Clinician Policy: The clinician policy directly reflects the observed actions taken by human clinicians in the MIMIC-IV dataset. It represents real-world practice and serves as an empirical reference point for evaluating the clinical plausibility of the learned policy.
  3. 3.
    Behavior Cloning (BC) Policy: The BC policy is a supervised learning model trained to imitate clinician decisions. Specifically, it is implemented as a neural network that minimizes the cross-entropy loss between the predicted actions and those recorded in the dataset. Although it may not replicate clinician behavior perfectly, its performance is expected to be similar to the clinician policy. This baseline helps evaluate whether the use of reinforcement learning yields meaningful improvements over simpler supervised learning approaches.
  4. 4.
    Expert Guidelines (EG) Policy: This rule-based policy encodes clinical recommendations drawn from the American Society for Parenteral and Enteral Nutrition (ASPEN) guidelines [compher2022guidelines], in combination with more recent evidence-based literature [wischmeyer2023personalized, stoian2025personalized], to reflect up-to-date, evidence-informed recommendations for enteral nutrition in critically ill patients. The policy is deterministic and constructed according to expert-derived nutritional targets and clinical rules. A full specification of this policy is provided in the Appendix (Section 3: Expert Guidelines Policy Definition).
3.2.3 Training and Hyperparameters
The patient trajectories were randomly split into 80% training and 20% testing sets. Our deep RL model is based on the D3QN framework coupled with CQL regularization. A grid search was conducted over CQL scaling coefficient $\alpha\in\{0.01,0.1,0.5,1\}$, and the discount factor $\gamma\in\{0.75,0.9,0.95,0.99\}$. For the network architecture, we evaluated different combinations of ReLU and sigmoid activation functions with 1-4 hidden layers, with layer widths drawn from $\{64,128,256,512\}$. All models were trained with a learning rate of 1e-4, batch size of 500.
The final model was selected based on offline evaluation performance, resulting the optimal hyperparameters to be $\alpha=0.5,\gamma=0.99$, ReLU activation on 3-layer network with decreasing dimensions (256, 128, and 64).
3.2.4 Off-Policy Evaluation
In online reinforcement learning (RL), policies are learned and evaluated via direct interaction with the environment. However, in the healthcare RL context—where the ”environment” corresponds to real patients—any real-time exploration is infeasible. Instead, we evaluate the performance of policies using various quantitative and qualitative off-policy evaluation (OPE) metrics tailored specifically to our target clinical outcomes. These do not require a simulator and instead adopt a well-established approach using offline learning on retrospective healthcare data.
  1. 1.
    Average Returns: We used Consistent Weighted Per-Decision Importance Sampling (CWPDIS) [thomas2016data] to estimate expected returns from each policy. As importance sampling based OPE methods require a behavior policy, we used our trained BC model baseline to estimate the behavior policy. A higher $V_{\text{CWPDIS}}$ corresponds to higher estimated returns and indicates a more effective policy.
  2. 2.
    Mortality vs. Estimated Returns: To evaluate the clinical alignment, We analyzed the correlation between estimated returns and observed mortality under the learned policies. An effective model should display a strong negative correlation, where higher expected returns translate to lower mortality. This indicates that the model learns to associate actions leading to lower returns with higher mortality, and vice versa.
  3. 3.
    Estimated Mortality Rate: We further computed the estimated mortality rates for the learned policies as follows. First, we categorized expected returns of patient trajectories into distinct bins and used mortality occurrences within each bin to calculate the average mortality rate. The resulting relationship between expected returns and mortality is then used to estimate mortality rates for all policies based on their expected returns.
  4. 4.
    Dosage Deviation vs. Mortality, and Biomarker Deviation: We qualitatively evaluated the performance of our learned policy by plotting the relationship between outcome metrics (mortality rate or next-state biomarker deviations) and dosage differences between clinician-chosen actions and policy-chosen actions. We estimated this relationship by categorizing quantile-level dosage differences into bins and computing mortality rates or deviation units for each bin. An effective policy should align with clinician dosages that resulted in positive outcomes ($x$-axis =0=0), and increasingly deviate from those associated with negative outcomes, ideally forming a U-shaped curve centered at 0. For biomarker deviation analysis, we focused on comparing calorie dosage differences against glucose deviations, and protein dosage differences against phosphate deviations for a clinically centered evaluation.
4 Results
We evaluated DeepEN against four baselines: clinician behavior, a random policy, behavior cloning (BC), and the expert guidelines policy. Our analysis spans both quantitative and qualitative dimensions, incorporating return-based metrics, estimated mortality, biomarker alignment, and policy behavior.
4.1 Quantitative Evaluation: Returns and Mortality
Table 2: Evaluation of Policies by Mortality Rate and CWPDIS
Policy Mortality Rate (%) CWPDIS
Clinician $22.5$ 5.87
Random $26.8\pm 0.49$ 1.62
BC $22.3\pm 0.15$ 5.95
EG $20.5\pm 0.29$ 8.11
DeepEN $\mathbf{18.8\pm 0.17}$ 11.89
DeepEN achieved the lowest estimated mortality among all tested policies, with a 3.7 ± 0.17 percent lower mortality rate compared to the clinician policy (18.8% vs. 22.5%) (Table 2). This corresponds to approximately 37 ± 1.7 fewer deaths per 1000 ICU patients. DeepEN also resulted the highest CWPDIS score among the policies, indicating superior expected returns under the learned value function.DeepEN also resulted the highest CWPDIS score among the policies, indicating superior expected returns under the learned value function (Table 2).
Note that, the guideline-based ASPEN policy shows improvement over clinicians, and DeepEN shows even further gains suggesting that offline RL can augment even expert-derived rules in enteral nutrition management.
4.2 Return-Mortality Correlation: Clinical Alignment
Figure 2: Mortality vs. expected returns. The shaded area represents the confidence interval.
Figure 2 displays a strong negative correlation between expected returns and mortality ($p<0.001$). This indicates that the policy learns to associate actions leading to lower returns with higher mortality and vice versa.
Figure 3: Dosage differences (x-axis) versus mortality (y-axis) for all policies. The shaded area indicates the confidence interval.
We analyzed how policy-recommended dosages differ from clinician choices, and their associated outcomes. Figure 3 shows that DeepEN achieves the closest approximation to the desired U-shaped relationship between dosage difference and mortality across all three components (calories, protein, and water). The lowest mortality is achieved when the policy recommendation closely matches the clinician’s choice (i.e., dosage difference $\approx 0$), and mortality increase with larger differences. This is evident that the DeepEN policy agrees with clinician behavior in favorable cases, and it deviates when clinician appears to recommend sub-optimal actions.
4.3 Dosage vs. Biomarker Deviations: Physiological Alignment
Figure 4: Dosage differences (x-axis) versus biomarker deviation (y-axis) for DeepEN. (A) Glucose deviation vs. calories dosage difference and (B) Phosphate deviation vs. protein dosage.
In Figure 4, we analyzed the relationship between dosage differences and next-state biomarker deviation to evaluate short-term physiological alignment. For both glucose and phosphate, DeepEN again showed the desired U-shaped pattern with minimal deviation when policy- and clinician- recommended dosages are similar.
This further validates that the intermediate biomarker rewards used during training were able to effectively guide the target toward actions that stabilize key metabolic indicators. Moreover, the model appears to effectively penalize both over- and under-nutrition, consistent with established physiological responses to EN.
5 Discussion and Conclusions
DeepEN applies deep RL to dynamically personalize enteral nutrition for critically ill patients, with the overarching goal of improving hospital survival. The framework is trained using an expert-informed state space and a physiologically grounded reward function that integrates long-term outcomes with short-term nutrition-relevant signals. Together, these design choices enable DeepEN to effectively distinguish actions associated with higher mortality from those linked to better survival, producing a policy that aligns with high-value clinician decisions while appropriately deviating from suboptimal ones. To ensure clinical plausibility, the framework incorporates a conservative Q-learning penalty that constrains the learned policy to remain within the support of observed clinical practice.
DeepEN significantly outperforms both clinician-derived and guideline-based policies across qualitative and quantitative evaluation metrics. The learned policy’s action distribution demonstrates closer adherence to established guideline recommendations than clinician practice, while still exhibiting meaningful and data-driven divergence (Appendix Section 4: Comparison of Action Distributions). Moreover, its superior performance relative to the behavior cloning policy indicates that the added complexity of reinforcement learning yields tangible benefits over simpler supervised learning approaches.
Although DeepEN’s observed mortality reduction of 3.7% (from 22.5% to 18.8%) is smaller than the improvements typically reported in healthcare RL studies focused on fluid and vasopressor dosage optimization [wu2023value, komorowski2018artificial], this effect remains clinically meaningful within the context of enteral nutrition therapy. The multifactorial nature of critical illness and the historically modest survival benefits demonstrated in large-scale nutrition trials [bear2017role, compher2022guidelines] underscore the significance of even incremental gains. Unlike hemodynamic interventions, which exert immediate effects on survival through cardiovascular stabilization, enteral nutrition confers its benefits more gradually through metabolic regulation, maintenance of gut integrity, and reduction of malnutrition-related complications. These mechanisms, though less acutely life-saving, are essential for supporting physiological recovery and long-term outcomes in critically ill patients. [adeyinka2018enteric]
Taken together, these findings demonstrate that personalized, reinforcement learning–based treatment strategies offer a distinct advantage over one-size-fits-all, guideline-based approaches, highlighting the potential of data-driven personalization to improve the safety and efficacy of nutritional care in the ICU.
While we train and evaluate DeepEN on a relatively large patient cohort, the data we use comes from a single hospital system. This may limit the generalizability of our model, as clinical practices and patient populations can vary substantially across institutions and regions. Validation on external cohorts will therefore be essential to confirm the robustness of our findings. However, the lack of detailed EN data in other large open-source critical care databases currently precludes this effort. We must also note that DeepEN’s reward function—while based heavily on well-established rewards functions in healthcare RL [raghu2017deep, tan2024advancing, wu2023value]—is handcrafted. Consequently, it may not optimally capture all relevant aspects of nutritional adequacy or clinical recovery. The exploration and development of alternative rewards functions that better integrate and capture short- and long-term goals relevant to EN, represents an important future direction.
Other promising directions for future research include the incorporation of multimodal data—particularly clinical text—to enhance contextual understanding of enteral nutrition (EN) decision making. Clinical notes can contain information critical to understanding feeding tolerance and complications, such as episodes of gastrointestinal (GI) intolerance (e.g., vomiting, diarrhea, abdominal distension) or risks of aspiration; this can be leveraged to improve the model’s ability to distinguish appropriate feeding interruptions from suboptimal underfeeding. Moreover, beyond dosage optimization, the model could be extended into a more holistic EN decision support tool capable of predicting optimal EN initiation timing, as well as nutrition-related complications such as refeeding syndrome, aspiration risk, or delayed gastric emptying.
6 Ethical Statement
In line with prior work on reinforcement learning–driven clinical decision support systems (CDSSs) for conditions such as sepsis [raghu2017deep], our proposed framework adopts a human-in-the-loop approach. The system is designed to assist, rather than replace, the role of clinicians by offering data-driven recommendations derived from past clinical trajectories. Importantly, the ultimate responsibility for treatment decisions remains with the clinician, preserving professional autonomy and ensuring that medical expertise governs final care plans. Our framework enables the integration of machine intelligence with clinical judgment, leveraging the strengths of both to improve patient outcomes without undermining ethical standards.
Additionally, our models were developed and validated solely on de-identified, publicly available retrospective data, with no involvement of real-time clinical interventions or patient contact. As such, the work poses no direct risk to patients and aligns with ethical research practices, minimizing concerns regarding safety, accountability, or consent in the current study phase.
7 Data Availability
The open-source MIMIC-IV 2.0 data used in this study is available at https://physionet.org/content/mimiciv/2.0/.
8 Code Availability
The underlying code for this study is available on Github via https://github.com/danjst/en_rl.
Acknowledgment
The authors would like to acknowledge the contributions of their team members and mentors from the Singapore IMAGINE AI 2024 Datathon.
Appendix
1 Cohort Information
Table 3: Cohort details
Cohort % Female Mean Age (years) Mean ICU Stay (hours) Total Population (n)
Overall 41.63 64.92 241 11378
Non-Survivors 43.19 68.76 258 2556
Survivors 41.18 63.81 237 8822
2 Action Information
Table 4: Distribution of Actions in Dataset
ID Cal Pro Water Count
0 1 1 1 21268
1 4 4 4 15932
2 4 4 3 13241
3 2 2 1 11755
4 4 4 2 9322
5 3 3 2 9179
6 1 1 2 9000
7 2 2 2 8683
8 1 1 4 8392
9 1 1 3 8093
10 3 3 3 7969
11 2 2 3 6763
12 3 3 1 5740
13 2 2 4 5536
14 3 3 4 5458
15 4 3 4 5169
16 3 4 2 4777
17 2 3 2 4695
18 3 4 3 4644
19 1 2 1 4325
20 4 3 3 4253
21 3 2 4 4098
22 2 3 1 3993
23 3 2 3 3485
24 2 1 4 $34$
25 3 4 4 3266
ID Cal Pro Water Count
26 3 2 2 3147
27 2 3 3 3128
28 4 3 2 2980
29 3 2 1 2352
30 4 4 1 2337
31 2 1 3 2232
32 1 2 2 1955
33 2 3 4 1774
34 2 1 2 1578
35 1 2 3 1384
36 4 3 1 1218
37 2 1 1 1205
38 3 4 1 1179
39 1 2 4 1062
40 4 2 4 1037
41 3 1 4 535
42 2 4 1 385
43 2 4 2 368
44 4 2 3 324
45 2 4 3 311
46 1 3 1 272
47 2 4 4 249
48 3 1 3 163
49 4 2 2 162
50 1 3 2 147
Table 5: Action Quantile Thresholds (4 hourly doses)
Quantile 1 2 3 4
Calories (kcal/kg) 0–1.91 1.91–3.05 3.05–4.13 >>4.13
Protein (g/kg) 0–0.08 0.08–0.14 0.14–0.19 >>0.19
Water (ml/kg) 0–3.61 3.61–5.40 5.40–8.15 >>8.15
3 Expert Guidelines Policy Definition
These nutritional targets are based on the 2021 ASPEN requirements for enteral nutrition [compher2022guidelines]. Original protein targets have been altered slightly with more recent evidence-based literature recommending a low-dosing period during the early-acute phase of critical illness [wischmeyer2023personalized, stoian2025personalized].
3.1 Calories
  • 1.
    For patients with BMI <30<30: provide 25 kcal/kg/day.
  • 2.
    For patients with BMI between $30$ and $50$: provide 22 kcal/kg/day.
  • 3.
    For patients with BMI >50>50: provide 11 kcal/kg/day.
  • 4.
    For all patients, only 70% of the above amounts are provided during the first 3 days of enteral feeding to simulate hypocaloric underfeeding.
3.2 Protein
3.2.1 Early Acute Phase (ICU Day 1-4)
  • 1.
    Day 1-2: 0.8 g/kg/day of protein.
  • 2.
    Day 3-4: 1.0 g/kg/day of protein.
3.2.2 Stable Phase (ICU Day >>4)
  • 1.
    For patients with BMI <30<30: provide 1.2 g/kg/day of protein.
  • 2.
    For patients with BMI between $30$ and $40$: provide 2 g/kg ideal body weight/day.
  • 3.
    For patients with BMI >40>40: provide 2.5 g/kg ideal body weight/day.
  • 4.
    Patients with burns receive 2 g/kg/day of protein, regardless of BMI.
  • 5.
    Patients on CRRT receive 2.5 g/kg/day of protein, regardless of BMI.
3.3 Water
  • 1.
    1.5 ml per 1 kcal of calories administered.
    4 Comparison of Action Distributions
    Figure 5: Action Distributions of a) Calories, b) Protein, and c) Water for Clinician, EG, and DeepEN policies.